# ITQ3_S: Interleaved Ternary Quantization with TurboQuant High-Fidelity 3-bit LLM Inference via Rotation-Domain Adaptive Quantization

### 30 Mar 2026 v1

@@ Abstract
Critically, we derive a mathematically rigorous dequantization procedure that inverts the FWHT exactly using a 256-point Inverse Walsh-Hadamard Transform fused into the CUDA shared-memory loading stage, ensuring zero-error round-trip fidelity between offline quantization and online inference. We prove that for any weight vector $\mathbf{w}\in\mathbb{R}^{256}$ processed by our pipeline, the reconstruction satisfies $\|\hat{\mathbf{w}}-\mathbf{w}\|_{2}\leq\epsilon_{q}$, where $\epsilon_{q}$ is determined solely by the ternary quantization grid and is strictly smaller than any uniform 3-bit baseline under equal bit-budget constraints.

### 31 Mar 2026 v2

@@ Abstract
Critically, we derive a mathematically rigorous dequantization procedure that inverts the FWHT exactly using a 256-point Inverse Walsh-Hadamard Transform fused into the CUDA shared-memory loading stage, ensuring that reconstruction error is bounded exclusively by the ternary quantization grid, with no additional error introduced by the transform inversion itself. We prove that for any weight vector $\mathbf{w}\in\mathbb{R}^{256}$ processed by our pipeline, the reconstruction satisfies $\|\hat{\mathbf{w}}-\mathbf{w}\|_{2}\leq\epsilon_{q}$, where $\epsilon_{q}$ is determined solely by the ternary quantization grid and is strictly smaller than that of uniform 3-bit baselines that do not exploit rotation-induced distribution normalization.
While TurboQuant establishes the theoretical foundation for FWHT-based rotation, it lacks a native CUDA kernel implementation, precluding direct deployment. Furthermore, naïvely composing TQ with existing weight quantizers — applying rotation only to KV cache while leaving weights in the original domain — introduces a systematic domain mismatch whose errors accumulate across transformer layers, ultimately degrading model quality below even standard 3-bit baselines. ITQ3_S resolves this by co-designing the FWHT rotation and the ternary quantization kernel as a single unified pipeline, grounding the rotation directly in the IQ3_S weight format and fusing the inverse transform into the CUDA MMQ kernel.